# Towards Scalable Automated Alignment of LLMs: A Survey

### v1

@@ 3.1.3 H3: LLMs can effectively learn in-context
From the perspective of automated alignment, ICL offers an efficient means to cold-start from a pre-trained LLM. With the assistance of ICL, just a few conversational samples in-context can yield a somewhat aligned model (Ganguli et al., 2022; Sun et al., 2023d; Lin et al., 2024a). Similarly, by prepending a few annotated exemplars in-context, ICL can also elicit judge and critic abilities of pre-trained LLM to some extent, or enhance the performance of original zero-shot capabilities (Bai et al., 2022b). Additionally, ICL presents a potential method for adaptive alignment (Xu et al., 2023c) to different social norms and regulations.
However, prepending few-shot exemplars in the context above can make inference inefficient (Gim et al., 2023) and interfered with the unrelated queries (Shi et al., 2023). Therefore, self-generated labels obtained from ICL could be directly used as pseudo labels, and distilled back into the LLMs only paired with the query. This paradigm is known as Context Distillation (Askell et al., 2021). For instance, in the alignment process of Llama-2 (Touvron et al., 2023), Context Distillation is used to alleviate the problem of long-term dependencies of system prompts. In Dromedary (Sun et al., 2023d), the base language model is transformed into a safe and helpful aligned model with minimal annotations by directly training on samples acquired from multiple ICL processes. Padmanabhan et al. (2023) demonstrate that Context Distillation can also be used to inject new knowledge to models by learning continuations from entity definitions; Furthermore, Yang et al. (2024b) illustrate the effectiveness of distilling the preference pair generated by contrastive in-context constraints back into the model.
@@
Unfortunately, the black box nature of ICL itself poses a significant challenge to alignment (Anwar et al., 2024). Without a comprehensive understanding of how LLMs learn in-context, the context distillation approach may introduce problems by potentially amplifying biases and errors inherent in the in-context learning process of the models. Moreover, the ability of long-term in-context learning (Agarwal et al., 2024) warrants further exploration, as it facilitates more efficient distillation and is crucial for scalable oversight settings where models need to comprehend lengthy professional documents or extensive self-play histories.
@@ 3.2.1 H4: Task decomposition
Task decomposition has long been regarded as an efficacious approach to tackling complex problems more effectively. For example, in cooperative games grounded in collective rationality, the overall benefits accrued by an alliance exceed the sum of individual gains (Shapley, 1971). Moreover, the divide-and-conquer paradigm and recursion are well-established and effective means employed in algorithm design for addressing problems of substantial scale and complexity (Hoare, 1961; Wilf, 2002).
@@
The discussion of this paradigm can be traced back to the assumption of factored cognition (Ought, 2017). It advocates that cognition tasks can be recursively decomposed. If an AI or human encounters a task that is difficult to solve, it can decompose the task, assign the decomposed problems to a series of its own copies for parallel processing, and finally merge these results. The copies focus on short-term work and work independently. A series of prompting methods implicitly or partially adopts the factored-cognition assumption for automated alignment. For instance, Zhou et al. (2023b) and Wang et al. (2023b) prompted the LLM to decompose the problem, then guide it to sequentially solve the sub-problems. It is also believed that task decomposition is an effective method for solving Easy-to-Hard generalization (Zhou et al., 2023b), that is, constructing decomposition prompts on simple samples and fill them in-context allows LLM the potential to generalize to difficult samples. Khot et al. (2023) further implemented recursive task decomposition.
@@
Based on the assumption of factored cognition, Iterative Distillation and Amplification (IDA) (Christiano et al., 2018) views each decomposition-merging process as a form of amplification and considers learning from the final merged results as a form of distillation. Although the original IDA paper builds this theoretical framework in a human-in-the-loop manner, where humans supervise the initial task decomposition step, it is likely that this process can be initiated without human oversight due to H1, H2, and H3 (Zhang and Parkes, 2023).
@@
Notably, Iterative Distillation and Amplification (IDA) represents a promising avenue towards achieving scalable oversight, making it possible to address long-horizon tasks that are difficult for humans to directly supervise by decomposing tasks into more tractable sub-problems. For example, labels for data points like “peer-review this survey” can take several months to collect in real world. Such problems can be tackled more quickly through factored cognition. Although some work partially demonstrates the effectiveness of IDA on real-world tasks like book-length summarization (Wu et al., 2021a), this ideology still relies on a set of crucial assumptions: 1) It is still unclear whether decompose a problem is the hardest part of solving it, if the cognitive burden cannot be distributed, IDA may struggle to take effect. 2) The error won’t accumulate. Although this paradigm does not require the collaboration between agents to be efficient (Christiano et al., 2018), too many errors can still be problematic. 3) The extent to which tasks can be parallelized. If the task-solving process is largely sequential, the time to collect signals might increase, but this appears to be a minor issue given the current deployment speed of LLMs. Overall, since these assumptions are difficult to prove or falsify, we advocate for more empirical research in this direction.
@@ 3.2.2 H5: Self-play
Complexity emerge from adversariality (Bansal et al., 2018). Self-play refers to a paradigm where an agent learns by iteratively playing games against itself, a form of non-cooperative game (Nash et al., 1951) in which each agent aims to maximize its own utility. Self-play serves as the foundation of many successful specialized superhuman AI systems like AlphaGo Zero (Silver et al., 2018) and StockFish (StockFish, 2023). Given this success, self-play seems to be a potential approach for enabling general proposed superhuman intelligence from LLMs. Two representative self-play methods are the Generator-Discriminator and the Debate approaches, with the latter involving $N\geq 2$ adversarial generators and one discriminator in a gaming environment.
@@
As discussed in H2, the judge and critic model is commonly considered a type of discriminator. For example, Yuan et al. (2024) utilized rewards from LLM-as-a-Judge to identify high-quality and low-quality responses from the generator, optimizing the generator towards the higher quality ones. However, the adversarial setting between the discriminator and the generator is limited because the only assumption is that the discriminator’s ability can be improved with general helpful training. The discriminator remains almost static (with unchanged prompts) during training, making it possible for the generator to be over-optimized against the discriminator, leading to reward hacking. Effectively improving the judge and critic module alongside the generator is thus a crucial problem. One reasonable approach is to use the online AIF setting, as suggested by Guo et al. (2024b). Another way to introduce a more adversarial setting is to optimize the gaming problem at inference time, as demonstrated in the Consensus Game by Jacob et al. (2024). It employs the piKL no-regret learning algorithm to iteratively update the strategies of both the generator and discriminator, converging to a Nash equilibrium. This equilibrium strategy is then used to rank candidate responses, prioritizing those agreed upon by both players.
As Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) have become a well-established class of methods in traditional NLP (Zhang et al., 2016; Wu et al., 2021b), another approach involves using a GAN-like discriminator to distinguish between the model’s current predicted distribution and the golden distribution. For instance, Chen et al. (2024g) found that a specific type of iterative DPO training, which consistently treats the policy-generated responses as negative and the golden responses as positive, can be viewed as a self-play process. In this process, the implicit reward of DPO functions as a discriminator between the model’s predictions and the golden samples. However, for open-ended questions, the golden distribution is sometimes still suboptimal, and this approach precludes the possibility of generating responses that are better than the golden ones.
@@
The debate paradigm (Irving et al., 2018) is largely inspired by the design of AlphaGo (Silver et al., 2018). In the learning algorithm of AlphaGo, three distinct components are integrated: a player, a counterpart—which is effectively a simulation of itself—and a value model that evaluates the win probability associated with each board state. By employing Monte Carlo Tree Search (MCTS), the algorithm conducts rollouts, defined as simulated self-play sequences that extend until a game’s conclusion. These rollouts enhance the accuracy of the value estimates through backward updates based on the outcomes, concurrently refining the policy by capitalizing on strategies that have previously led to victories.
@@
This provide a possible oversight solution to build trustworthy superhuman AI systems. Irving et al. (2018) show that the honesty is the better strategy than lie in debate paradigm through proof-of-concept experiments. As a extension of this, Brown-Cohen et al. (2023) propose a new set of debate protocols, wherein the honest strategy can always succeed through a simulation involving only a polynomial number of steps. Khan et al. (2024) conduct a thorough empirical study on the feasibility of implementing the debate paradigm on LLMs: it was found that the debate paradigm can significantly enhance truthfulness, and more persuasive (Anthropic, 2024) debaters lead to more truthful outcomes.
This offers a potential solution to construct trustworthy superhuman AI systems. Irving et al. (2018) honesty proves to be a superior strategy compared to deceit in a debate paradigm through proof-of-concept experiments. Expanding on this notion, Brown-Cohen et al. (2023) introduced a novel set of debate protocols, wherein the honest approach can consistently prevail, as demonstrated in simulations requiring only a polynomial number of steps. Khan et al. (2024) undertook a comprehensive empirical investigation into the practicality of implementing the debate paradigm on LLMs, revealing that it substantially enhances truthfulness, with more compelling debaters resulting in more truthful outcomes (Anthropic, 2024).
Apart from the classic natural language debate, a series of works have implemented the debate paradigm in game scenarios. A representative arena is the bargaining task (Nash et al., 1950). Fu et al. (2023a) focused on zero-sum variants of bargaining, where the balloon seller aims to sell at a higher price while the buyer seeks a lower price. They observed significant variations in bargaining capabilities among different LLMs and their capacity to learn from play experiences and feedback. Cheng et al. (2024) implemented the adversarial language game adversarial taboo (Yao et al., 2021), where an attacker and a defender engage in a conversation centered around a target word visible only to the attacker. The attacker subtly induces the defender to unconsciously utter the target word, while the defender tries to avoid doing so and guess the word from the context. Both players acquire basic gaming skills through imitation learning from a teacher LLM and then refine their strategies through self-play. Interestingly, less capable player LLMs not only improve their win rates in this specific game through self-play but also enhance their general reasoning abilities Ma et al. (2023a) introduced the red-teaming game, a more intricate adversarial team game where LLMs are initialized as a joint set of red-teaming policies to prompt the target LLM to produce harmful content. They proposed a solver to ensure the final meta-strategy approximates a Nash equilibrium within a certain $\epsilon$ margin.
@@ 3.2.3 Discussion
Task decomposition and self-play both necessitate that LLMs function as agents. However, the challenge of aligning LLMs as agents is more complex compared to aligning them as chatbots, as it requires considerations of behavior-level alignment (Pan et al., 2023), the dynamics of environment and self-constraints (Garrabrant and Demski, 2018; Shavit et al., 2023; Yang et al., 2024d). We emphasize the significance of this research direction and advocate for increased efforts in this domain. Furthermore, an even more intricate problem lies in proving the theoretical safety and trustworthiness of multi-agent systems. Although research in this area is nascent (Yang and Wang, 2020; DiGiovanni and Zell, 2021), advancements in game theory (Hazra and Anjaria, 2022), automated theorem proving techniques (Polu and Sutskever, 2020) and real-world simulation technology (Brooks et al., 2024) may offer insights into addressing this challenge.
While adversarial self-play offers a potential solution to the scalability oversight challenge, it can also give rise to the emergence of more persuasive and autonomous agents (Tao et al., 2024). Such developments could have significant social impacts and ethical risks, such as the potential for models to generate more persuasive articles than humans, which could be exploited for political manipulation. Encouragingly, several prominent model engine providers have taken steps to monitor and mitigate these potential side effects. For instance,OpenAI’s Preparedness team has established benchmarks related to persuasion and autonomy (OpenAI, 2023a). They categorized model risks into four levels: low, medium, high, and critical, stipulating that models above the high-risk threshold cannot be developed, and those above medium risk cannot be deployed. We call upon the academic community and third-party organizations to dedicate additional efforts to the development and review of a robust safety framework for highly capable agents.
@@ 4.1 Instruction Construction
Collecting large-scale instructions with high quality and diversity serves as the foundation for achieving alignment through behavior imitation. The most intuitive strategy involves filtering out high-quality data from human-written instructions, such as those found in ShareGPT  [^3] https://huggingface.co/datasets/anon8231489123/ShareGPT_Vicuna_unfiltered. However, this approach requires substantial human effort and expertise, which also introduces significant noise. Consequently, many studies focus on utilizing LLMs for automatic instruction generation, thereby significantly reducing the dependence on human annotation. Based on the information provided for instruction construction, there are currently three representative strategies:
In-Context Generation, which provides in-context demonstrations to guide LLMs generating instructions. For example, Unnatural Instructions (Honovich et al., 2023) and Self-Instruct (Wang et al., 2023f) begin with a small set of human-written instructions. These instructions are randomly selected to create context examples that prompt LLMs to generate additional instructions. Similarly, Stanford Alpaca (Taori et al., 2023) employs a comparable approach by using ChatGPT to gather instruction-response pairs.
However, the initial seed instructions may lack diversity, which limits the range of generated instructions. To address this, LaMini-LM (Wu et al., 2024) not only utilizes gpt-3.5-turbo for Example-Guided Instruction Generation based on In-context Learning (ICL), but also incorporates wiki data for Topic-Guided Instruction Generation, thereby constructing a large, offline distilled instruction dataset. Guo et al. (2024a) use high-quality samples relevant to the target domain as examples for ICL to generate additional samples. Dynosaur (Yin et al., 2023) employs ChatGPT to generate instructions by leveraging meta-information from existing NLP datasets, creating a dynamically growing instruction tuning dataset. Furthermore, LLM2LLM (Lee et al., 2024) enhances the dataset by generating new data from the teacher model using incorrect predictions from the student model, which improves the dataset’s ability to handle data-constrained tasks.
@@
Instruction Evolution, which involves the process of rewriting existing instructions based on predefined handwritten evolution principles. Xu et al. (2024a) apply an instruction evolution paradigm, where existing instruction data is transformed using randomly selected handwritten evolution principles. This approach is employed for robust models such as GPT-3, facilitating the rewriting of instructions, thereby reducing the need for manual annotation and enhancing the model’s ability to manage complex tasks. Building upon the Evol-Instruct framework, TeaMs-RL (Gu et al., 2024) advances this methodology by training another model through reinforcement learning to generate optimized instruction trajectories.
@@
Instruction Backtranslation, which employs LLMs to predict instructions that correspond to the golden responses extracted from human handwritten text or web documents. Humpback (Li et al., 2024e) and LongForm (Köksal et al., 2024) utilize corresponding prompts to label text from web documents, thereby creating high-quality instructional data. Differently, TEGIT (Chen et al., 2023e) engages ChatGPT to annotate sampled text, simultaneously generating instructions, inputs, and outputs to effectively filter out noise. REInstruct (Chen et al., 2024b) builds high-quality instructions from an unlabelled corpus and rewrites the unlabelled text to enhance its quality as response.

### v2

@@ 3.1.3 H3: LLMs can effectively learn in-context
From the perspective of automated alignment, ICL offers an efficient means to cold-start from a pre-trained LLM. With the assistance of ICL, just a few conversational samples in-context can yield a somewhat aligned model (Ganguli et al., 2022; Sun et al., 2023d; Lin et al., 2024a). Similarly, by prepending a few annotated exemplars in-context, ICL can also elicit judge and critic abilities of pre-trained LLM to some extent, or enhance the performance compared to the zero-shot setting (Bai et al., 2022b). Additionally, ICL presents a potential method for adaptive alignment (Xu et al., 2023c) to different social norms and regulations.
However, prepending few-shot exemplars in the context above can make inference inefficient (Gim et al., 2023) and interfered with the unrelated queries (Shi et al., 2023). Therefore, self-generated labels obtained from ICL could be directly used as pseudo labels, and distilled back into the LLMs only paired with the query. This paradigm is known as Context Distillation (Askell et al., 2021). For instance, in the alignment process of Llama-2 (Touvron et al., 2023), Context Distillation is used to alleviate the problem of long-term dependencies of system prompts. In Dromedary (Sun et al., 2023d), the base language model is transformed into a safe and helpful aligned model with minimal annotations by directly training on samples acquired from multiple ICL processes. Padmanabhan et al. (2023) demonstrate that Context Distillation can also be used to inject new knowledge to models by learning continuations from entity definitions. Furthermore, Yang et al. (2024b) illustrate the effectiveness of distilling the preference pair generated by contrastive in-context constraints back into the model.
@@
Unfortunately, the black box nature of ICL itself poses a significant challenge to alignment (Anwar et al., 2024). Without a comprehensive understanding of how LLMs learn in-context, the context distillation approach may introduce problems by potentially amplifying biases and errors inherent in the ICL process of the models. Moreover, the ability of long in-context learning (Agarwal et al., 2024; Li et al., 2024e) warrants further exploration, as it facilitates more efficient distillation and is crucial for scalable oversight settings where models need to comprehend lengthy professional documents or extensive self-play histories.
@@ 3.2.1 H4: Task decomposition
Task decomposition has long been regarded as an effective approach to tackling complex problems (Lee and Anderson, 2001). For example, in cooperative games grounded in collective rationality, the overall benefits accrued by an alliance exceed the sum of individual gains (Shapley, 1971). Moreover, the divide-and-conquer paradigm and recursion are well-established and effective means employed in algorithm design for addressing problems of substantial scale and complexity (Hoare, 1961; Wilf, 2002).
@@
The discussion of this paradigm can be traced back to the assumption of factored cognition (Ought, 2017). It advocates that cognition tasks can be recursively decomposed. If an AI or human encounters a task that is difficult to solve, it can decompose the task, assign the decomposed problems to a series of its own copies for parallel processing, and finally merge these results. The copies focus on short-term work and work independently. A series of prompting methods implicitly or partially adopts the factored-cognition assumption for automated alignment. For instance, Zhou et al. (2023b) and Wang et al. (2023b) prompt the LLM to decompose the problem, then guide it to sequentially solve the sub-problems. It is also believed that task decomposition is an effective method for solving Easy-to-Hard generalization (Zhou et al., 2023b), that is, constructing decomposition prompts on simple samples and filling them in-context allows LLM the potential to generalize to difficult samples. Khot et al. (2023) further implement recursive task decomposition.
@@
Based on the assumption of factored cognition, Iterative Distillation and Amplification (IDA) (Christiano et al., 2018) views each decomposition-merging process as a form of amplification and considers learning from the final merged results as a form of distillation. Although the original IDA paper builds this theoretical framework in a human-in-the-loop manner, where humans supervise the initial task decomposition step, it is likely that this process can be initiated without much human oversight given H1, H2, and H3 (Zhang and Parkes, 2023).
@@
Notably, IDA represents a promising avenue towards achieving scalable oversight, making it possible to address long-horizon tasks that are difficult for humans to directly supervise, by decomposing tasks into more tractable sub-problems. For example, labels for data points like “peer-review this survey” can take several months to collect in real world. Such problems can be tackled more quickly through factored cognition. Although some work partially demonstrates the effectiveness of IDA on real-world tasks like book-length summarization (Wu et al., 2021a) and complex code bug fixing (Wen et al., 2024), this ideology still relies on a set of crucial assumptions: 1) It is still unclear whether decompose a problem is the hardest part of solving it, if the cognitive burden cannot be distributed, IDA may struggle to take effect. 2) The error won’t accumulate. Although this paradigm does not require the collaboration between agents to be efficient (Christiano et al., 2018), too many errors can still be problematic. 3) The extent to which tasks can be parallelized. If the task-solving process is largely sequential, the time to collect signals might increase, but this appears to be a minor issue given the current deployment speed of LLMs. Overall, since these assumptions are difficult to prove or falsify, we advocate for more empirical research in this direction.
@@ 3.2.2 H5: Self-play
Complexity emerges from adversariality (Bansal et al., 2018). Self-play refers to a paradigm where an agent learns by iteratively playing games against itself, a form of non-cooperative game (Nash et al., 1951) in which each agent aims to maximize its own utility. It serves as the foundation of many successful specialized superhuman AI systems like AlphaGo Zero (Silver et al., 2018) and StockFish (StockFish, 2023). Given this success, self-play seems to be a potential approach for enabling general proposed superhuman intelligence from LLMs. Two representative self-play methods are the Generator-Discriminator and the Debate approaches, with the latter involving $N\geq 2$ adversarial generators and one discriminator in a gaming environment.
@@
As discussed in H2, the judge and critic model is commonly considered a type of discriminator. For example, Yuan et al. (2024b) utilized rewards from LLM-as-a-Judge to identify high-quality and low-quality responses from the generator, optimizing the generator towards the higher quality ones. However, the adversarial setting between the discriminator and the generator is limited because the only assumption is that the discriminator’s ability can be improved with general helpful training. The discriminator remains almost static (the prompts are unchanged) during training, making it possible for the generator to be over-optimized against the discriminator, leading to reward hacking. Effectively improving the judge and critic module alongside the generator is thus a crucial problem. One reasonable approach is to use the online AIF setting, as suggested by Guo et al. (2024a). Moreover, Cheng et al. (2023) propose an adversarial game between the policy and reward model, facilitated through a constrained min-max loss. Another way to introduce a more adversarial setting is to optimize the gaming problem at inference time, as demonstrated in the Consensus Game by Jacob et al. (2024). It employs the piKL no-regret learning algorithm to iteratively update the strategies of both the generator and discriminator, converging to a Nash equilibrium. This equilibrium strategy is then used to rank candidate responses, prioritizing those agreed upon by both players.
As Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) have become a well-established class of methods in traditional NLP (Zhang et al., 2016; Wu et al., 2021b), another line of work involves using a GAN-like discriminator to distinguish between the model’s current predicted distribution and the golden distribution. For instance, Chen et al. (2024g) find that a specific type of iterative DPO training, which consistently treats the policy-generated responses as negative and the golden responses as positive, can be viewed as a self-play process. In this process, the implicit reward function of DPO serves as a discriminator between the model’s predictions and the golden samples. Expanding on this, Shaikh et al. (2024) further add replay comparison signals between earlier iteration of models and the golden, and comparisons between a model and its successive model in the self-play process. However, for open-ended questions, the golden distribution is sometimes still suboptimal, and this setting precludes the possibility of generating responses that are better than the golden ones.
@@
The debate paradigm (Irving et al., 2018) is largely inspired by factored cognition and AlphaGo (Silver et al., 2018). In the learning algorithm of AlphaGo, three distinct components are integrated: a player, a counterpart (itself), and a value model that evaluates the win rate associated with each board state. By employing Monte Carlo Tree Search (MCTS), the algorithm conducts rollouts, which are simulated self-play sequences that extend until a game’s conclusion. These rollouts enhance the accuracy of the value estimates through backward updates based on the outcomes, concurrently refining the policy by capitalizing on strategies that have previously led to victories.
@@
This provide a possible oversight solution to build trustworthy superhuman AI systems. Irving et al. (2018) show that honesty is better strategy than lie in debate paradigm through proof-of-concept experiments. As an extension of this, Brown-Cohen et al. (2023) propose a new set of debate protocols, wherein the honest strategy can always succeed through a simulation involving only a polynomial number of steps. Khan et al. (2024) conduct a thorough empirical study on the feasibility of implementing the debate paradigm on LLMs: it was found that the debate paradigm can significantly enhance truthfulness, and more persuasive (Anthropic, 2024a) debaters lead to more truthful outcomes.
Apart from the classic natural language debate, a series of works have implemented the debate paradigm in game scenarios. A representative arena is the bargaining task (Nash et al., 1950). Fu et al. (2023a) focus on zero-sum variants of bargaining, where the balloon seller aims to sell at a higher price while the buyer seeks a lower price. They observed significant variations in bargaining capabilities among different LLMs and their capacity to learn from play experiences and feedback. Cheng et al. (2024) implement the adversarial language game adversarial taboo (Yao et al., 2021), where an attacker and a defender engage in a conversation centered around a target word visible only to the attacker. The attacker subtly induces the defender to unconsciously utter the target word, while the defender tries to avoid doing so and guess the word from the context. Both players acquire basic gaming skills through imitation learning from a teacher LLM and then refine their strategies through self-play. Interestingly, less capable player LLMs not only improve their win rates in this specific game but also enhance their general reasoning abilities. Ma et al. (2023a) introduce the red-teaming game, a more intricate adversarial team game where LLMs are initialized as a joint set of red-teaming policies to prompt the target LLM to produce harmful content. They propose a solver to ensure the final meta-strategy approximates a Nash equilibrium within a certain $\epsilon$ margin. Zheng et al. (2024c) suggests addressing the alignment problem by allowing an attacker to prompt a defender LLM to generate answers that might result in low rewards, while the defender tries to maximize the rewards of these prompts. The solution of this game is considered an iterative min-max optimization process with constraints.
@@ 3.2.3 Discussion
Task decomposition and self-play both necessitate LLMs functioning as agents. However, the challenge of aligning LLMs as agents is more complex compared to aligning them as chatbots, as it requires considerations of behavior-level alignment (Pan et al., 2023), the dynamics of environment and self-constraints (Garrabrant and Demski, 2018; Shavit et al., 2023; Yang et al., 2024f). We emphasize the significance of this research direction and advocate for increased efforts in this domain. Furthermore, an even more intricate problem lies in proving the theoretical safety and trustworthiness of multi-agent systems. Although research in this area is nascent (Yang and Wang, 2020; DiGiovanni and Zell, 2021), advancements in game theory (Hazra and Anjaria, 2022), automated theorem proving techniques (Polu and Sutskever, 2020) and real-world simulation technology (Brooks et al., 2024) may offer insights into addressing this challenge.
While adversarial self-play offers a potential solution to the scalable oversight challenge, it can also give rise to the emergence of more deceptive (Hubinger et al., 2024), persuasive and autonomous agents (Tao et al., 2024). Such developments could have significant social impacts and ethical risks, such as the potential for models to generate more persuasive articles than humans, which could be exploited for political manipulation. Encouragingly, several prominent model engine providers have taken steps to monitor and mitigate these potential side effects. For instance, OpenAI’s Preparedness team has established benchmarks related to persuasion and autonomy (OpenAI, 2023a). They categorize model risks into four levels: low, medium, high, and critical, stipulating that models above the high-risk threshold cannot be developed, and those above medium risk cannot be deployed. We call upon the academic community and third-party organizations to dedicate additional efforts to the development and review of a robust safety framework for highly capable agents.
@@ 4.1 Instruction Construction
Collecting large-scale instructions with high quality and diversity serves as the foundation for achieving alignment through behavior imitation. The most intuitive strategy involves filtering out high-quality data from human-written instructions. However, this approach requires substantial human effort and expertise, which also introduces significant noise. Consequently, many studies focus on utilizing LLMs for automatic instruction generation, thereby significantly reducing the dependence on human annotation. Based on the information provided for instruction construction, there are currently 3 representative strategies:
In-Context Generation, which provides in-context demonstrations to guide LLMs generating instructions. For example,  Honovich et al. (2023); Wang et al. (2023f); Taori et al. (2023) begin with a small set of human-written instructions. These instructions are randomly selected to create context examples that prompt LLMs to generate additional instructions. To further improve the scale and diversity of generated instructions, LaMini-LM (Wu et al., 2024a) additionally introduces wiki data for topic-guided instruction generation, thereby constructing a large, offline distilled instruction dataset. Dynosaur (Yin et al., 2023) leverages meta-information from existing NLP datasets to create a dynamically growing instruction tuning dataset. Moreover, LLM2LLM (Lee et al., 2024) enhances the difficulty and complexity of instructions by iteratively introducing examples that the model fails to answer correctly.
@@
Instruction Evolution, which involves rewriting existing instructions based on pre-defined evolution principles. Evol-Instruct (Xu et al., 2024a) employ LLMs to conduct instruction evolution based on handwritten principles, thereby reducing the need for manual annotation and enhancing the model’s ability to manage complex tasks. Building upon this, TeaMs-RL (Gu et al., 2024) trains another model through reinforcement learning to generate optimized evolution trajectories. Considering the reliance on manually written principles, Auto Evol-Instruct (Zeng et al., 2024b) proposes an automated principle construction method, further enhancing the diversity and complexity of evolved instructions.
@@
Instruction Backtranslation, which employs LLMs to predict instructions based on responses extracted from human handwritten text or web documents. LongForm (Köksal et al., 2024), TEGIT (Chen et al., 2023e) and Humpback (Li et al., 2024f) prompt LLM to construct instructions according to cleaned web corpus. REInstruct (Chen et al., 2024b) builds instructions from an unlabelled corpus and rewrites the unlabelled text to enhance its quality as a response.